If two charged balloons are 20cm apart and they feel a force of electrical repulsion of 50N, what would the force of electrical repulsion become if the balloons were moved apart to a distance of 100cm?
1
Expert's answer
2020-05-25T10:55:52-0400

According to Coulomb's law, the force of electrical repulsion is

"F = k_e\dfrac{q_1q_2}{r^2}" . Therefore, if we increase the distance 5 times, the force decreases "5^2=25" times. We may write it as

"\dfrac{F_2}{F_1} = \dfrac{k_e\dfrac{q_1q_2}{r_2^2}}{k_e\dfrac{q_1q_2}{r_1^2}} = \dfrac{r_1^2}{r_2^2} = \left(\dfrac{r_1}{r_2} \right)^2 = \left(\dfrac{20}{100} \right)^2 = \dfrac{1}{25}."

So "F_2 = F_1/25= 50\,\mathrm{N}/25 = 2\,\mathrm{N}."